OpenStudy (anonymous):

Consider the sequence 2, 6, 18, 54, . . . Let "n" = the term number in the sequence. Let A(n) = the value of the "n"th term of the sequence. Complete each statement: a. A(1) = 2 = 2 * 3^?

OpenStudy (anonymous):

0

OpenStudy (campbell_st):

the general equation is A(n) = 2*3^(n -1) n >=0

OpenStudy (anonymous):

The formula is Un=U1(r)^n-1 The part you're looking for is the n-1 as n is 1, the equation must be raised to 0.

OpenStudy (campbell_st):

so A(1) = 2*3^1 A(2) = 2*3^1 etc A(3) = 2*3^2